OpenStudy (anonymous):

HELP PLEASE FAN AND MEDAL A number cube with faces labeled 1–6 is rolled 72 times. About how many times would the result be 4? A. 4 B. 6 C. 8 D. 12

OpenStudy (anonymous):

@Destinymasha

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

i think it is 12

OpenStudy (anonymous):

@johnweldon1993 HELP

OpenStudy (anonymous):

can u check my answer please i said 12

OpenStudy (loser66):

yup

OpenStudy (loser66):

Again, Yup, you are correct

OpenStudy (anonymous):

wait

OpenStudy (johnweldon1993):

Yeah 12 is correct....since the probability will be \(\large \frac{1}{6}\) each time....so \[\large \frac{1}{6}\times 72 = \frac{72}{6} = 12\]

OpenStudy (anonymous):

can u help with more